Evaluate the the integral of (2x+1)^7 dx from-½ to ½


1
Expert's answer
2021-02-23T13:01:25-0500

"\begin{array}{l}
\int\limits_{ - \frac{1}{2}}^{\frac{1}{2}} {{{\left( {2x + 1} \right)}^7}} dx = \frac{1}{2}\int\limits_{ - \frac{1}{2}}^{\frac{1}{2}} {2{{\left( {2x + 1} \right)}^7}} dx = \frac{1}{2}\int\limits_{ - \frac{1}{2}}^{\frac{1}{2}} {{{\left( {2x + 1} \right)}^7}} d(2x + 1) = \left. {\frac{{{{\left( {2x + 1} \right)}^8}}}{{16}}} \right|_{ - \frac{1}{2}}^{\frac{1}{2}} = \\
 = \frac{{{{\left( {1 + 1} \right)}^8} - {{( - 1 + 1)}^8}}}{{16}} = 16
\end{array}"

Answer: 16